Find the maximum and minimum values of inductance that can be obtained

using ten 10mH inductors.


1
Expert's answer
2021-12-06T12:21:16-0500

 Series inductances


"L_{TC}=L_1+L_2+...+L_{10}"

"=10(0.01H)=0.1H"

Parallel  inductances


"\dfrac{1}{L_{TP}}=\dfrac{1}{L_1}+\dfrac{1}{L_2}+...+\dfrac{1}{L_{10}}"

"=10(\dfrac{1}{0.01H})"

"L_{TP}=\dfrac{0.01H}{10}=0.001H=1mH"

The maximum possible inductance is "0.1H=100mH."

The mimimum possible inductance is "0.001H=1mH."